reasonableness: Which of the following order pairs is not included on the graph equation y=x+9
(7,16) (15,24)(20,29) (28,36)

Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20055)

y=x+9
(7,16) (15,24)(20,29) (28,36)
Let's check to see if (7,16) is included on the graph of y-x+9.  
The first coordinate of (7,16) is 7. That's the x-coordinate, 
so that's what we substitute for x.  So instead of the x in y=x+9 
we will substitute 7.

The second coordinate of (7,16) is 16. That's the y-coordinate, 
so that's what we substitute for y.  So instead of the y in y=x+9 
we will substitute 16.

So 

y=x+9

becomes

16=7+9

and when we add 7+9 we get 16, so we end up with

16=16

So (7,16) is included on the graph of y=x+9

----------------------------

Next let's check to see if (15,24) is included on the graph of y-x+9.  
The first coordinate of (15,24) is 15. That's the x-coordinate, 
so that's what we substitute for x.  So instead of the x in y=x+9 
we will substitute 15.

The second coordinate of (15,24) is 24. That's the y-coordinate, 
so that's what we substitute for y.  So instead of the y in y=x+9 
we will substitute 24.

So 

y=x+9

becomes

24=15+9

and when we add 15+9 we get 24, so we end up with

24=24

So (15,24) is included on the graph of y=x+9

----------------------------

----------------------------

Next let's check to see if (20,29) is included on the graph of y-x+9.  
The first coordinate of (20,29) is 20. That's the x-coordinate, 
so that's what we substitute for x.  So instead of the x in y=x+9 
we will substitute 20.

The second coordinate of (20,29) is 29. That's the y-coordinate, 
so that's what we substitute for y.  So instead of the y in y=x+9 
we will substitute 29.

So 

y=x+9

becomes

29=20+9

and when we add 20+9 we get 29, so we end up with

29=29

So (20,29) is included on the graph of y=x+9

---------------------------

Finally let's check to see if (28,36) is included on the graph of y-x+9.  
The first coordinate of (28,36) is 28. That's the x-coordinate, 
so that's what we substitute for x.  So instead of the x in y=x+9 
we will substitute 28.

The second coordinate of (28,36) is 36. That's the y-coordinate, 
so that's what we substitute for y.  So instead of the y in y=x+9 
we will substitute 36.

So 

y=x+9

becomes

36=28+9

and when we add 28+9 we get 37, so we end up with

36=37

That's FALSE!  So (28,36) is NOT included on the graph of y=x+9

So the answer is (28,36).

Here's what we have shown about the graph.  The black line below
is the graph of y=x+9, and you see that all the ordered pairs given
represent points that are included in the line graph except (28,36),
which is not on the line.



Edwin

Answer by stanbon(75887)

Which of the following order pairs is not included on the graph equation y=x+9
(7,16) (15,24)(20,29) (28,36)
---
Answer: (28,36)
====================
Cheers,
Stan H.
=============